# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v1

@@ Appendix 0.B More Implementation Details
0.B.2 Continual learning pipeline.
Algorithm 1 Continual learning pipeline
0:  The whole representation for dynamic scenes.
[…]
2:  while $N_{chunk}$ do
[…]
5:     if $k~{}\mathbf{mod}~{}T_{episode}~{}=0$  then
[…]
7:     else
[…]
10:     end if
11:     Train the $k^{th}$ model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
[…]
13:  end while
Appendix 0.C More Details about the Long Video Dataset
We collect the dataset for evaluating the scalability of our method and further research. We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes. The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes nor affect the anonymity of the submission. All videos are collected with the explicit consent of all the participants. The view in the upper center is selected for validation. We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses.
Appendix 0.D More Ablation Studies
0.D.1 Temporal representations.
Our continual learning pipeline is also compatible with a variety of representations for the time axis. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with them on DyNeRF dataset and show the results in Tab. 9. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
@@ Appendix 0.D More Ablation Studies
Table 9: Ablation study on field composition methods.
@@
0.D.2 Field composition methods.
We also conduct ablation studies on the field composition methods and show the results in Tabs. 9 and 9. The field composition process is defined as:
@@
Full composition in Tab. 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Tab. 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
@@
0.D.3 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tabs. 11 and 11. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. Our method is robust within a wide range of chunk sizes. As in Tab. 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Tab. 11. We do not cherry-pick the best chunk size and select $T_{chunk}=10$ as the default configuration in our method. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.

### v3

@@ 3.4 Ablation studies
3.4.3 Temporal representations.
Our continual learning pipeline is also compatible with a variety of representations for the time axis. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with the ones mentioned above on DyNeRF dataset and show the results in Table 8. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
@@
Table 9: Ablation study on the field composition methods.
@@
3.4.4 Field composition methods.
We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 10. The field composition process is defined as:
@@
Full composition in Table 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 10, the compared composition methods in Figs. 10(c) and 10(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
@@
3.4.5 Regularization weights.
We also provide ablation studies on the regularization weights of auxiliary spatial encodings. We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Table 10. The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$.
Table 10: Ablation study on regularization weight $\lambda_{r}$.
[…]
$10^{-5}$ 30.20 0.0269 0.198
$10^{-4}$ 30.18 0.0268 0.196
$10^{-3}$ 30.23 0.0269 0.198
$10^{-2}$ 30.09 0.0279 0.197
1 29.66 0.0313 0.205
3.4.6 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tables 11 and 12. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. Our method is robust within a wide range of chunk sizes. As in Table 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 12. We select $T_{chunk}=10$ as the default configuration. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.
@@
4 Conclusion and Limitation
Appendix A Detailed results on the DyNeRF dataset